OpenStudy (anonymous):

(6-(4.3+1.2)) If I could please have some help. I can't tell if it's a negative or minus sign.:)

OpenStudy (freckles):

6-(4.3+1.2) ?

OpenStudy (freckles):

perform the operations in the grouping symbol

OpenStudy (anonymous):

you first do 4.3 + 1.2

OpenStudy (freckles):

4.3+1.2=?

OpenStudy (freckles):

oops kevin beat me

OpenStudy (anonymous):

The whole equation looks like this > 8.6 * (6 -(4.3+1.2))but I don't understand what to do with the (-)

OpenStudy (anonymous):

it a system of elimination

OpenStudy (anonymous):

Kevin, I understand that I first have to to solve what is in the parenthesis but what does the (-) do?

OpenStudy (anonymous):

it means minus

OpenStudy (anonymous):

(6-(4.3+1.2)) (6-(5.5)) = (6 - 5.5)

OpenStudy (anonymous):

6-5.5 = what?

OpenStudy (anonymous):

Wow!!! FACE PALM -.-! Thanks:))

OpenStudy (anonymous):

lol! your welcome friend.